OpenStudy (anonymous):

**I give medals** Which of the following ordered pairs is the solution to the system of linear equations? (5, 14) (4, 4) (2, 5) (6, 3) x plus two y equals twelve; y equals three x minus one

OpenStudy (igreen):

x + 2y = 12 y = 3x - 1 Just plug in the values and see which one fits both equations. Let's try (2, 5) 2 + 2(5) = 12 2 + 10 = 12 12 = 12 So the first equation fits, now let's try the 2nd one: y = 3x - 1 5 = 3(2) - 1 5 = 6 - 1 Can you subtract 6 - 1? @kyrea

OpenStudy (anonymous):

thx